OpenStudy (anonymous):

Find an equation of the line that goes through the point (5,4) and has a slope 0f 2. Leave equation in point-slope form.

OpenStudy (anonymous):

point slope form: y-y1 = m(x-x1) where m is slope, and (x1, y1) is a point that the line goes through plug in what you know and youve got an answer

OpenStudy (anonymous):

@etemplin has answered your question.

OpenStudy (anonymous):

Is it y-4=m(x-5)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

For m, you plug in the slope

OpenStudy (anonymous):

you can substitute 2 for m because 2 is the slope and m is the slope

OpenStudy (anonymous):

You stated that the slope is 2, so just plug that in.

OpenStudy (anonymous):

y-4=2(x-5)

OpenStudy (anonymous):

Yep